Shift report 20011108 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Nov 08 2001 - 23:56:18 EST

  • Next message: Ian Bearden: "Run plan feed back..."

    ------------------------------------------------------------
    Date: Thu Nov  8 23:56:18 2001, Shift: 16:00-24:00
    Supervisor: rw, with: 
    Log book#10 pages 116 - 116
    
    Shift Summary:
    ------------------------------------------------------------
    1. Scheduled maintenance finished at 20:30. At 22:20, 
    "preparing to inject into RHIC" was announced. Several 
    atempts were made to store the beam into RHIC, but no 
    successful storage (not ramping!) into RHIC was made.  
    
    2. none
    
    3.No new problems. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Nov 08 2001 - 23:56:40 EST